OpenStudy (abbles):

Explain how the graph of each given function is a transformation of the graph of y = x^2

OpenStudy (abbles):

y = (-3x)^2

OpenStudy (abbles):

I'm having a bit of trouble with this lesson. Horizontal and vertical stretches and whatnot.

OpenStudy (prathamesh_m):

Is it \[-3x ^{2} or (-3x)^{2}\]

OpenStudy (abbles):

According to my lesson, y = c*f(x) stretches the graph of f(x) vertically by a factor of c when c>1. y = c*f(x) shrinks the graph of f(x) vertically to c times as tall when c<1. y = f(cx) shrinks the graph of f(x) horizontally to 1/c times as wide when c>1. y = f(x/c) stretches the graph of f(x) horizontally by a factor of c when c>1. I don't see a formula for y = f(cx) when c<1, which this seems to be.

OpenStudy (abbles):

It is (-3x)^2, the second one.

OpenStudy (prathamesh_m):

Your question would then simply be \[y=9x ^{2}\] Since \[(-3x)^{2} = 9x ^{2}\]

OpenStudy (prathamesh_m):

So in this case c would be 9

OpenStudy (abbles):

Ah. Interesting

OpenStudy (abbles):

You're sure?

OpenStudy (prathamesh_m):

yeah i am

OpenStudy (abbles):

So the answer would be "The graph y = x^2 would stretch vertically by a factor of 9" or something along those lines?

OpenStudy (abbles):

Mind clarifying a few more @Prathamesh_M ?

OpenStudy (abbles):

I need to do the same thing with y = x^ - 5 and y = (x - 5)^2

OpenStudy (prathamesh_m):

Second one is simple. For \[y=x ^{2}\] The vertex is the origin. For \[y=(x-5)^{2}\] the only change is that the vertex shifts to (5,0)

OpenStudy (prathamesh_m):

I'm not sure about the first one though

OpenStudy (abbles):

So it moves 5 to the right? for the first one?

OpenStudy (abbles):

What would the difference between y = x^2 - 5 and y = (x - 5)^2 be?

OpenStudy (prathamesh_m):

For \[y=(x-5)^{2}\]

OpenStudy (prathamesh_m):

\[y=x ^{2}-5\] is the same as \[y+5=x ^{2}\]

OpenStudy (prathamesh_m):

which means that the vertex would be (0,-5) instead of (5,0). The plots are identical in all other aspects

OpenStudy (abbles):

Ah okay... so the first equation would move down 5 units? and the second would move right 5?

OpenStudy (prathamesh_m):

yeah exactly

OpenStudy (prathamesh_m):

Basically you can express all such parabolas in the form of \[(y-y _{1})=4a(x-x _{1})^{2}\]

OpenStudy (abbles):

Thank you so much! I just have one more left that I'm really confused about. y = -2x^2

OpenStudy (prathamesh_m):

For which the vertex would be (x1,y1)

OpenStudy (abbles):

Is that vertex form?

OpenStudy (prathamesh_m):

yes

OpenStudy (prathamesh_m):

That last one would be a downward open parabola

OpenStudy (prathamesh_m):

so you flip it upside down and stretch it by a factor of two

OpenStudy (prathamesh_m):

'it' here refers to \[y=x ^{2}\]

OpenStudy (abbles):

In the lesson I was given this: y = c*f(x) shrinks the graph of f(x) vertically to c times as tall when c<1. Would it mean the same thing to say: "The graph y = x^2 would shrink vertically to 2 times as tall" ? I don't really understand that.

OpenStudy (prathamesh_m):

Do u mean for that last one?

OpenStudy (abbles):

Yes

OpenStudy (abbles):

Because y = -2x^2 seems to fit that criteria.

OpenStudy (prathamesh_m):

The graph actually shrinks only when modulus of c is less than one

OpenStudy (prathamesh_m):

when u see a negative sign it means its downward open

OpenStudy (abbles):

Isn't -2 less than one?

OpenStudy (prathamesh_m):

yeah but i said modulus of -2

OpenStudy (prathamesh_m):

which would be 2

OpenStudy (prathamesh_m):

so actually what u typed before is wrong it should have modulus of 'c'

OpenStudy (abbles):

Are you saying this: "The graph y = x^2 would shrink vertically to 2 times as tall" should be -2 instead of 2?

OpenStudy (prathamesh_m):

so graph shrinks only when c is something like 1/5

OpenStudy (abbles):

Oh I gotcha..

OpenStudy (prathamesh_m):

it would then shrink by a factor of 5

OpenStudy (abbles):

Got it. That makes sense. So for this problem, The graph would stretch vertically (or horizontally?) by a factor of 2 and also open downward?

OpenStudy (abbles):

Would that also make it a vertical reflection?

OpenStudy (abbles):

@Prathamesh_M

OpenStudy (prathamesh_m):

yup it would look like a water image of the graph stretched by a factor of 2.